# Business Informatics Research Center, Cluj-Napoca, Romania

### v1

@@ Abstract
Recent work shown the capability of Large Language Models (LLMs) to solve tasks related to Knowledge Graphs, such as Knowledge Graph Completion, even in Zero- or Few-Shot paradigms. However, they are known to hallucinate answers, or output results in a non-deterministic manner, thus leading to wrongly reasoned responses, even if they satisfy the user’s demands. To highlight opportunities and challenges in knowledge graphs-related tasks, we experiment with two distinguished LLMs, namely Mixtral-8x7B-Instruct-v0.1, and gpt-3.5-turbo-0125, on Knowledge Graph Completion for static knowledge graphs, using prompts constructed following the TELeR taxonomy, in Zero- and One-Shot contexts, on a Task-Oriented Dialogue system use case. When evaluated using both strict and flexible metrics measurement manners, our results show that LLMs could be fit for such a task if prompts encapsulate sufficient information and relevant examples.
@@
Knowledge Graphs (KGs) are defined as graphs of data intended to accumulate and convey knowledge of the real world [3]. Their nodes represent entities of interest and edges represent potentially different relations between these entities.
KGs are integrated into various systems to enhance their abilities of storing and processing information. In our previous work [4], we focused on developing an ontology-enhanced Task-Oriented Dialogue (TOD) system equipped with a static KG capable of mapping the context of the discussion and storing relevant information. Numerous benefits stem from adopting this approach, including enabling concurrent threads of conversation within a single discourse and utilizing the KG to validate data as a proxy. Additionally, the system gains the capability to execute Create-Retrieve-Update-Delete (CRUD) operations on domain-specific KGs. Specifically, we employed the Knowledge Graph Completion (KGC) task to construct two KGs and the Knowledge Graph Reasoning (KGR) task to handle CRUD operations. KGC’s objective is to deduce absent information within a specified KG [8], drawing from either input text or pre-existing knowledge. Nonetheless, our system relies on input text template-matching rules, constraining the authenticity of dialogues and hindering adaptability to novel concepts beyond the predefined ontology. While embedding deep learning models into the TOD system architecture demonstrated encouraging outcomes, we still failed to completely address the previously mentioned limitations.
@@
Our experiments explore LLMs for static KGs contexts. Two well-established LLMs are used: Mixtral-8x7B-Instruct-v0.1111https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1 [6], and gpt-3.5-turbo-0125222https://platform.openai.com/docs/models/gpt-3-5-turbo, each possessing different properties. Communicating with such models involves the use of prompts, which are natural language instructions formatted in such way that the model understands the user’s intent. We test their capabilities of solving the aforementioned task using multiple prompting styles, including human-created and model-specific rephrased ones. Each prompt belongs to a level defined according to the TELeR taxonomy [9], that includes techniques such as Direct Prompting (DP), In-Context Learning (ICL), or Chain of Thought (COT), under Zero- and One-Shot contexts. To illustrate an appropriate application scenario for LLMs and KG tasks, we extract sample phrases from the training phase of our TOD system. Two datasets are obtained, including one with an increased difficulty, with test cases requiring reasoning steps that are not explicitly mentioned in the prompts. This approach allows us to not only evaluate the capability of LLMs in addressing the KG-specific tasks, but also to investigate their synergy with TOD systems. Finally, we report the accuracy and F1 scores of each LLM on both dataset, under strict and flexible measurement paradigms.
@@ 2 Related Work
As opposed to the above mentioned literature, we increase the number of textual inputs, expanding the generality of our conclusions. Similar to Khorashadizadeh et al. [7], we test the capacity of a proprietary LLM - namely gpt-3.5-turbo-0125 on the KGC task. Moreover, we include an open-source LLM - Mixtral-8x7B-Instruct-v0.1 [6], to facilitate research on open-source models, given their greater adaptability and cost-effectiveness compared to proprietary alternatives. To the best of our knowledge, we are among the first to test Mixtral for KG-related tasks. Another difference from them is that our prompts are more diverse and easier to track, as they are leveled according to the TELeR taxonomy [9]. We introduce flexible metrics for gauging additional post-processing efforts. Finally, we also test the possibility of integrating an LLM with an ontology-enhanced TOD system, to sharpen its natural language processing and KG-related capabilities, by utilizing sample phrases from its training routine, resulting in two datasets, differentiated by their level of difficulty.
@@ 3.1 Datasets format and distribution
Table 1 presents examples from each category. Details about the explicit types are available in the project’s repository. Each text is accompanied by its related relationships, values, intent, and text type. Using regex templates, the information is converted into a dictionary. At the end, we obtained two datasets: Templates Easy (TE) and Templates Hard (TH). The first dataset includes explicit and misleading text types, while the second one benefits from the addition of implicit-type texts. Table 2 presents the distribution of texts per class type, on each dataset.
@@ 3.2 Prompt engineering
To better encapsulate information and facilitate the replication of experiments, we utilize regex templates to convert each dictionary object - representing a text from a dataset, into a Prompt object. Such objects hold many details, such as the system prompt, its version and level, the input text, its type and mentioned class type, the golden labels and the ones that could be accepted alternatively, the system message order, and metadata information about each model’s prediction of the input prompt. For a more comprehensive analysis, we enable the adjustment of the system’s message position within the final prompt. Practice suggests333For prompt engineering we also followed OpenAI (https://community.openai.com) and HuggingFace (https://huggingface.co/docs/transformers/main/tasks/prompting) suggestions that positioning the system’s prompt after the user’s message could potentially enhance the performance of LLMs by mitigating long-context memory limitations. Lastly, we provide the option to flatten each Prompt object into dictionaries that are placed in text files for future reuse.
@@ 3.3 Metrics
The flexible paradigm proposed here allows formatting mistakes that can be corrected in post-processing steps, or triples that are partially true to be counted as being accurate. This flexible measurement allows one to positively evaluate models that might not be such precise, but require fewer resources than more efficient ones.
Our flexible measurement paradigm modifies the strict calculation of the accuracy and F1 scores by including certain penalty for each instance that could be considered valid, even if it does not exactly follow the given prompt. Below, we list the penalties we considered for each sort of LLM output errors.
@@
Another category of penalties is related to the information content of a triple. For example, let’s ask a model to construct a simple ID for each given instance of a class - specifically, the capitalized name of the class concatenated with “1”. We have noticed that some models tend to replace the number “1” with another single digit. Thus, if altering the final digit of a predicted identifier to “1” signifies correctness for a triple, the model is subjected to a penalty of .
@@ 4 Results and discussion
Experiments were conducted on Google Colab, utilizing a virtual machine equipped with two Intel Xeon CPU 2.20GHz processors. We experimented with Mixtral-8x7B-Instruct-v0.1 and gpt-3.5-turbo-0125. Mixtral is open-source, leveraging the Mixture of Experts[6] architecture, consisting of eight sub-networks, each of 7B parameters, accounting for a total of 56B parameters. GPT-3.5 Turbo is a well-known proprietary model that represents a fine-tuned version of GPT 3, consisting of 175B parameters. For Mixtral-8x7B-Instruct-v0.1, we used the HuggingFace Serverless API endpoint, whereas for gpt-3.5-turbo-0125 queries were directed to OpenAI’s official API.
Each experiment was iterated three times, involving 4500 prompts in total, with each run lasting approximately 90 minutes. Interaction with Mixtral consumed about of the experimentation time. The cost incurred by OpenAI’s model totaled around 5USD, while the HuggingFace endpoint generated no cost. Each set of predictions could be loaded, tested and visualized from the paper’s repository, available at https://github.com/IonutIga/LLMs-for-KGC.
@@
Model | Mixtral-8x7B-Instruct-v0.1 | gpt-3.5-turbo-0125 | Total | |
Level | Metric | strict | flexible | strict | flexible | strict | flexible |
1 | Accuracy | 0.23 | 0.47 | 0.34 | 0.45 | 0.28 | 0.46 |
F1 | 0.25 | 0.58 | 0.41 | 0.58 | 0.33 | 0.58 | |
2 | Accuracy | 0.19 | 0.49 | 0.41 | 0.51 | 0.30 | 0.50 |
F1 | 0.18 | 0.53 | 0.47 | 0.61 | 0.33 | 0.58 | |
3 | Accuracy | 0.19 | 0.44 | 0.30 | 0.42 | 0.24 | 0.43 |
F1 | 0.20 | 0.55 | 0.38 | 0.58 | 0.29 | 0.57 | |
4.1 | Accuracy | 0.25 | 0.63 | 0.88 | 0.88 | 0.56 | 0.76 |
F1 | 0.25 | 0.63 | 0.88 | 0.88 | 0.56 | 0.76 | |
4.2 | Accuracy | 0.19 | 0.69 | 0.85 | 0.87 | 0.52 | 0.78 |
F1 | 0.19 | 0.75 | 0.85 | 0.87 | 0.52 | 0.81 | |
Total Accuracy | 0.21 | 0.54 | 0.55 | 0.62 | 0.38 | 0.58 | |
Total F1 | 0.22 | 0.61 | 0.60 | 0.72 | 0.41 | 0.67 |
Model | Mixtral-8x7B-Instruct-v0.1 | gpt-3.5-turbo-0125 | Total | |
Level | Metric | strict | flexible | strict | flexible | strict | flexible |
1 | Accuracy | 0.38 | 0.50 | 0.39 | 0.46 | 0.38 | 0.48 |
F1 | 0.42 | 0.59 | 0.47 | 0.57 | 0.44 | 0.58 | |
2 | Accuracy | 0.15 | 0.37 | 0.43 | 0.46 | 0.29 | 0.41 |
F1 | 0.17 | 0.49 | 0.51 | 0.58 | 0.34 | 0.53 | |
3 | Accuracy | 0.20 | 0.42 | 0.39 | 0.46 | 0.29 | 0.44 |
F1 | 0.22 | 0.51 | 0.49 | 0.59 | 0.35 | 0.56 | |
4.1 | Accuracy | 0.19 | 0.58 | 0.85 | 0.89 | 0.52 | 0.73 |
F1 | 0.19 | 0.59 | 0.85 | 0.89 | 0.52 | 0.73 | |
4.2 | Accuracy | 0.42 | 0.74 | 0.84 | 0.89 | 0.63 | 0.81 |
F1 | 0.42 | 0.78 | 0.84 | 0.89 | 0.63 | 0.83 | |
Total Accuracy | 0.27 | 0.52 | 0.58 | 0.63 | 0.43 | 0.58 | |
Total F1 | 0.28 | 0.59 | 0.64 | 0.71 | 0.46 | 0.65 |
Model | Mixtral-8x7B-Instruct-v0.1 | gpt-3.5-turbo-0125 | Total | |
Level | Metric | strict | flexible | strict | flexible | strict | flexible |
1 | Accuracy | 0.25 | 0.41 | 0.34 | 0.42 | 0.30 | 0.42 |
F1 | 0.28 | 0.48 | 0.41 | 0.53 | 0.35 | 0.51 | |
2 | Accuracy | 0.08 | 0.30 | 0.43 | 0.50 | 0.26 | 0.39 |
F1 | 0.08 | 0.35 | 0.49 | 0.59 | 0.29 | 0.47 | |
3 | Accuracy | 0.09 | 0.35 | 0.28 | 0.40 | 0.19 | 0.38 |
F1 | 0.10 | 0.46 | 0.35 | 0.52 | 0.23 | 0.50 | |
4.1 | Accuracy | 0.15 | 0.47 | 0.77 | 0.77 | 0.46 | 0.62 |
F1 | 0.15 | 0.49 | 0.77 | 0.78 | 0.46 | 0.63 | |
4.2 | Accuracy | 0.00 | 0.47 | 0.75 | 0.76 | 0.37 | 0.62 |
F1 | 0.00 | 0.56 | 0.75 | 0.76 | 0.37 | 0.66 | |
Total Accuracy | 0.11 | 0.39 | 0.51| 0.57 | 0.31 | 0.48 | |
Total F1 | 0.12 | 0.47 | 0.56 | 0.65 | 0.34 | 0.66 |
Model | Mixtral-8x7B-Instruct-v0.1 | gpt-3.5-turbo-0125 | Total | |
Level | Metric | strict | flexible | strict | flexible | strict | flexible |
1 | Accuracy | 0.33 | 0.43 | 0.32 | 0.41 | 0.32 | 0.42 |
F1 | 0.37 | 0.50 | 0.38 | 0.52 | 0.37 | 0.51 | |
2 | Accuracy | 0.14 | 0.35 | 0.39 | 0.46 | 0.26 | 0.41 |
F1 | 0.15 | 0.45 | 0.45 | 0.56 | 0.30 | 0.58 | |
3 | Accuracy | 0.12 | 0.39 | 0.38 | 0.45 | 0.25 | 0.42 |
F1 | 0.13 | 0.46 | 0.47 | 0.55 | 0.30 | 0.51 | |
4.1 | Accuracy | 0.07 | 0.47 | 0.71 | 0.75 | 0.39 | 0.60 |
F1 | 0.07 | 0.48 | 0.71 | 0.75 | 0.39 | 0.61 | |
4.2 | Accuracy | 0.31 | 0.56 | 0.70 | 0.74 | 0.50 | 0.65 |
F1 | 0.31 | 0.61 | 0.70 | 0.74 | 0.50 | 0.68 | |
Total Accuracy | 0.19 | 0.44 | 0.50 | 0.56 | 0.35 | 0.55 | |
Total F1 | 0.20 | 0.50 | 0.55 | 0.63 | 0.38 | 0.57 |
@@
Elaborate instructions do not necessarily yield better results. Upon analyzing both types of prompts across all levels, it appears that augmenting the prompt with more information does not consistently enhance performance. Level 3 prompts, when evaluated rigorously, exhibit an average decline of in accuracy and in F1 score compared to levels 1 and 2. However, when evaluated using more flexible metrics, the discrepancy diminishes to an average of only for accuracy and for F1 score. This variance could be attributed to the inclusion of explanatory text by language models, particularly by Mixtral 8x7B, as they strive to replicate the input text.
@@
ICL and COT prompting techniques lead to best results. Both models scored their best results when prompted at levels 4.1 and 4.2, no matter the dataset or prompting template. It is no surprise that such models work best when an adequate output example is given, as literature [9] suggests. However, as Mixtral 8x7B sometimes provides explanations for its output, erroneous reasoning steps are noticeable, especially in cases where the input text contains a class type that is not present in the ontology. Thus, despite gpt-3.5-turbo-0125 exhibiting this behavior less frequently, LLMs still have significant room for improvement in terms of reasoning capabilities.
@@
Mixtral 8x7B rarely follows the required output format. The two metric measurement paradigms offer valuable insights into the models capacity to follow the given prompts. While GPT-3.5 Turbo exhibits minimal disparity between the two perspectives, Mixtral 8x7B rarely produces texts that align with the specified template. Common errors include the addition of explanatory text (as evidenced by the 0 scores at the 4.2 level in table 6) or the full IRI of an entity. When strictly evaluated, the open-source model only tops accuracy on the Template Easy (TE), while on flexible paradigm it reaches accuracy on the same dataset. Comparatively, the proprietary model outputs accuracy under both metrics measurements.
@@
Asking models to rephrase the system prompt is generally a good idea. Some experiments in the literature [8] ask the LLMs to formulate prompts for a given task. Inspired by it, we ask the LLMs to rephrase our manually written prompts to better align with their capabilities. As a comparison, Mixtral 8x7B benefits the most under rigorous evaluation, with an average increase of in accuracy and F1 score. GPT-3.5 Turbo seems to conserve its behavior, signaling an increase of only . Nonetheless, it’s promising to see the open-source model enhancing its output by closely adhering to the provided system prompt.

### v2

@@ Abstract
Recent work has shown the capability of Large Language Models (LLMs) to solve tasks related to Knowledge Graphs, such as Knowledge Graph Completion, even in Zero- or Few-Shot paradigms. However, they are known to hallucinate answers, or output results in a non-deterministic manner, thus leading to wrongly reasoned responses, even if they satisfy the user’s demands. To highlight opportunities and challenges in knowledge graphs-related tasks, we experiment with three distinguished LLMs, namely Mixtral-8x7b-Instruct-v0.1, GPT-3.5-Turbo-0125 and GPT-4o, on Knowledge Graph Completion for static knowledge graphs, using prompts constructed following the TELeR taxonomy, in Zero- and One-Shot contexts, on a Task-Oriented Dialogue system use case. When evaluated using both strict and flexible metrics measurement manners, our results show that LLMs could be fit for such a task if prompts encapsulate sufficient information and relevant examples.
@@
Knowledge Graphs (KGs) are defined as graphs of data intended to accumulate and convey knowledge of the real world [4]. Their nodes represent entities of interest and edges represent potentially different relations between these entities. KGs are integrated into various systems to enhance their abilities of storing and processing information.
Task-Oriented Dialogue (TOD) systems, alongside chatbots, are conversational agents possessing capabilities of engaging in natural language dialogues with human users. Different from chatbots, TOD systems aim to solve the user’s specific tasks within certain domains [1]. In our previous work [6], we focused on developing an ontology-enhanced TOD system equipped with a static KG capable of mapping the context of the discussion and storing relevant information. Numerous benefits stem from adopting this approach, including enabling concurrent threads of conversation within a single discourse and utilizing the KG to validate data as a proxy. Additionally, the system gains the capability to execute Create-Retrieve-Update-Delete (CRUD) operations on domain-specific KGs. The acronym CRUD refers to the four basic operations that can be executed against persistent storages, such as relational or object databases, or other types of knowledge bases like KGs, to create, maintain or update them. As TOD systems aim to solve a variety of tasks, we decided that enabling such basic, but important ones is a suitable use-case to start with. Specifically, we employed the Knowledge Graph Completion (KGC) task to construct two KGs and the Knowledge Graph Reasoning (KGR) task to handle CRUD operations. KGC’s objective is to deduce absent information within a specified KG [10], drawing from either input text or pre-existing knowledge.
Nonetheless, our TOD system relies on input text template-matching rules, constraining the authenticity of dialogues and hindering adaptability to novel concepts beyond the predefined ontology. Hence, in a subsequent study [5], we experimented with training neural networks - specifically fine-tuning BERT, a pre-trained model, to discern user intent and pertinent associated entities from input text. While embedding deep learning models into the TOD system architecture demonstrated encouraging outcomes, we still failed to completely address the previously mentioned limitations.
@@
Our experiments explore LLMs for static KGs contexts. Two well-established LLMs are used: Mixtral-8x7b-instruct-v0.1111https://huggingface.co/mistralai/Mixtral-8x7b-Instruct-v0.1 [8], and GPT-3.5-Turbo-0125222https://platform.openai.com/docs/models/gpt-3-5-turbo, alongside the newest GPT model, the GPT-4o333https://platform.openai.com/docs/models/gpt-4o version, each possessing different properties. Communicating with such models involves the use of prompts, which are natural language instructions formatted in such way that the model understands the user’s intent. We test their capabilities of solving the aforementioned task using multiple prompting styles, including human-created and model-specific rephrased ones. Each prompt belongs to a level defined according to the TELeR taxonomy [11], that includes techniques such as Direct Prompting (DP), In-Context Learning (ICL), or Chain of Thought (COT), under Zero- and One-Shot contexts. To illustrate an appropriate application scenario for LLMs and KG tasks, we extract sample phrases from the training phase of our TOD system. Two datasets are obtained, including one with an increased difficulty, with test cases requiring reasoning steps that are not explicitly mentioned in the prompts. This approach allows us to not only evaluate the capability of LLMs in addressing the KG-specific tasks, but also to investigate their synergy with TOD systems. Finally, we report the accuracy and F1 scores of each LLM on both dataset, under strict and flexible measurement paradigms.
@@ 2 Related Work
As opposed to the above mentioned literature, we increase the number of textual inputs, expanding the generality of our conclusions. Similar to Khorashadizadeh et al. [9], we test the capacity of a proprietary LLM - namely GPT, with two versions: GPT-3.5-Turbo-0125 and GPT-4o on the KGC task. Moreover, we include an open-source LLM - Mixtral-8x7b-Instruct-v0.1 [8], to facilitate research on open-source models, given their greater adaptability and cost-effectiveness compared to proprietary alternatives. To the best of our knowledge, we are among the first to test Mixtral for KG-related tasks. Another difference from them is that our prompts are more diverse and easier to track, as they are leveled according to the TELeR taxonomy [11]. We introduce flexible metrics for gauging additional post-processing efforts. Finally, we also test the possibility of integrating an LLM with an ontology-enhanced TOD system, to sharpen its natural language processing and KG-related capabilities, by utilizing sample phrases from its training routine, resulting in two datasets, differentiated by their level of difficulty.
@@ 3.1 Datasets format and distribution
Table 1 presents examples from each category. Details about the explicit types are available in the project’s repository444https://github.com/IonutIga/LLMs-for-KGC. Each text is accompanied by its related relationships, values, intent, and text type. Using regex templates, the information is converted into a dictionary. At the end, we obtained two datasets: Templates Easy (TE) and Templates Hard (TH). The first dataset includes explicit and misleading text types, while the second one benefits from the addition of implicit-type texts. Table 2 presents the distribution of texts per class type, on each dataset.
@@ 3.2 Prompt engineering
To better encapsulate information and facilitate the replication of experiments, we utilize regex templates to convert each dictionary object - representing a text from a dataset, into a Prompt object. Such objects hold many details, such as the system prompt, its version and level, the input text, its type and mentioned class type, the golden labels and the ones that could be accepted alternatively, the system message order, and metadata information about each model’s prediction of the input prompt. For a more comprehensive analysis, we enable the adjustment of the system’s message position within the final prompt. Practice suggests555For prompt engineering we also followed OpenAI (https://community.openai.com) and HuggingFace (https://huggingface.co/docs/transformers/main/tasks/prompting) suggestions that positioning the system’s prompt after the user’s message could potentially enhance the performance of LLMs by mitigating long-context memory limitations. Lastly, we provide the option to flatten each Prompt object into dictionaries that are placed in text files for future reuse.
Three important paradigms are usually employed when designing prompts [14]: Zero-, One- and Few-Shot. The former’s prompt includes only the objective’s description and the input data that should be processed. One-Shot includes exactly one example of how the task should be solved against some different input data. Intuitively, Few-Shot refers to multiple, relevant examples added to the prompt. After selecting the paradigm, one should decide about the prompting technique [14]. The current work employes three main approaches, as follows: Direct Prompting which refers to a prompt that only comprises the task’s description and the input to work on; In-Context Learning (ICL) which adds a relevant example of a solution to the given task on a different input data, and Chain of Thought (COT) which expands the prompt with the exemplified solution’s intermediary reasoning steps.
@@ 3.3 Metrics
The flexible paradigm proposed here allows formatting mistakes that can be corrected in post-processing steps, or triples that are partially true to be counted as being accurate. This flexible measurement allows one to positively evaluate models that might not be such precise, but require fewer resources than more elaborate ones.
Our flexible measurement paradigm modifies the strict calculation of the accuracy and F1 scores by including certain penalty for each instance that could be considered valid, even if it does not exactly follow the given prompt. In cases of erroneous formatted output, deductions are computed from the accuracy (100%) of the entire prediction for each input text, whereas inaccuracies in the triple content lead to deductions only concerning the accuracy of a certain fact. Below, we list the penalties we considered for each sort of LLM output errors.
@@
Another category of penalties is related to the information content of a triple. For example, let’s ask a model to construct a simple ID for each given instance of a class - specifically, the capitalized name of the class concatenated with “1”. We have noticed that some models tend to replace the number “1” with another single digit. Thus, if altering the final digit of a predicted identifier to “1” signifies correctness for a triple, the model is subjected to a penalty of . This percentage value adheres to the three-component structure of a fact, such that, if one part is wrong, while the other two are correct, the model should still gain benefit of its prediction. This method of evaluation only applies to the validity of a constructed ID(s). Any other type of mistake is not allowed, since it alters the factuality of the implied information.
The penalty values described above and considered in our paper fit our specific task. We emphasize that these values are not fixed, and someone who wants to adapt the flexible measurement paradigm for another task is free to change them, according with her specific needs.
@@ 4 Results and discussion
Experiments were conducted on Google Colab, utilizing a virtual machine equipped with two Intel Xeon CPU 2.20GHz processors. We experimented with Mixtral-8x7b-Instruct-v0.1, GPT-3.5-Turbo-0125 and GPT-4o. Mixtral is open-source, leveraging the Mixture of Experts[8] architecture, consisting of eight sub-networks, each of 7B parameters, accounting for a total of 56B parameters. GPT-3.5-Turbo-0125 is a well-known proprietary model that represents a fine-tuned version of GPT 3, consisting of 175B parameters. GPT-4o boasts over 200B parameters, being the latest OpenAI model, advertised as their best performer. For Mixtral-8x7b-Instruct-v0.1, we used the HuggingFace Serverless API endpoint, whereas for GPT-3.5-Turbo-0125 and GPT-4o queries were directed to OpenAI’s official API.
Each experiment was iterated three times, involving aprox. 6750 prompts in total, with each run lasting approximately 120 minutes. Interaction with Mixtral consumed about of the experimentation time. GPT-4o generated a cost around 40USD, while the GPT-3.5-Turbo-0125 only about 5USD. For Mixtral, the HuggingFace endpoint generated no cost. Each set of predictions could be loaded, tested and visualized from the paper’s repository, available at https://github.com/IonutIga/LLMs-for-KGC.
We notice that for the GPT models, an extra post-processing step is required, after receiving the produced output. Due to their ability to generate JSON formatted output, it surrounds its response with a specific tag (i.e. "“‘json…“‘"). One solution is to include a guideline in the prompt to avoid this behavior, but very rarely, around of times, it still adds it. Thus, to ensure that prompts are identical for all models, and be sure that the tag is not present in the output, we post-process the GPT output in our code. We do this to enable a fair analysis solely of the output text.
@@
Model | Mixtral | GPT-3.5-Turbo | GPT-4o | Total | |
Level | Metric | strict | flexible | strict | flexible | strict | flexible | strict | flexible |
1 | Accuracy | 0.23 | 0.47 | 0.34 | 0.45 | 0.61 | 0.63 | 0.39 | 0.52 |
F1 | 0.25 | 0.58 | 0.41 | 0.58 | 0.73 | 0.75 | 0.46 | 0.64 | |
2 | Accuracy | 0.19 | 0.49 | 0.41 | 0.51 | 0.63 | 0.85 | 0.41 | 0.62 |
F1 | 0.18 | 0.53 | 0.47 | 0.61 | 0.64 | 0.89 | 0.43 | 0.68 | |
3 | Accuracy | 0.19 | 0.44 | 0.30 | 0.42 | 0.71 | 0.86 | 0.40 | 0.57 |
F1 | 0.20 | 0.55 | 0.38 | 0.58 | 0.72 | 0.89 | 0.43 | 0.67 | |
4.1 | Accuracy | 0.25 | 0.63 | 0.88 | 0.88 | 0.89 | 0.91 | 0.67 | 0.81 |
F1 | 0.25 | 0.63 | 0.88 | 0.88 | 0.89 | 0.91 | 0.67 | 0.81 | |
4.2 | Accuracy | 0.19 | 0.69 | 0.85 | 0.87 | 0.89 | 0.91 | 0.64 | 0.82 |
F1 | 0.19 | 0.75 | 0.85 | 0.87 | 0.89 | 0.91 | 0.64 | 0.84 | |
Total Accuracy | 0.21 | 0.54 | 0.55 | 0.62 | 0.75 | 0.83 | 0.50 | 0.67 | |
Total F1 | 0.22 | 0.61 | 0.60 | 0.72 | 0.78 | 0.87 | 0.53 | 0.73 |
Model | Mixtral | GPT-3.5-Turbo | GPT-4o | Total | |
Level | Metric | strict | flexible | strict | flexible | strict | flexible | strict | flexible |
1 | Accuracy | 0.38 | 0.50 | 0.39 | 0.46 | 0.62 | 0.64 | 0.46 | 0.53 |
F1 | 0.42 | 0.59 | 0.47 | 0.57 | 0.73 | 0.75 | 0.54 | 0.64 | |
2 | Accuracy | 0.15 | 0.37 | 0.43 | 0.46 | 0.36 | 0.85 | 0.31 | 0.56 |
F1 | 0.17 | 0.49 | 0.51 | 0.58 | 0.36 | 0.90 | 0.35 | 0.66 | |
3 | Accuracy | 0.20 | 0.42 | 0.39 | 0.46 | 0.07 | 0.77 | 0.22 | 0.55 |
F1 | 0.22 | 0.51 | 0.49 | 0.59 | 0.07 | 0.73 | 0.26 | 0.61 | |
4.1 | Accuracy | 0.19 | 0.58 | 0.85 | 0.89 | 0.66 | 0.90 | 0.57 | 0.79 |
F1 | 0.19 | 0.59 | 0.85 | 0.89 | 0.66 | 0.88 | 0.59 | 0.79 | |
4.2 | Accuracy | 0.42 | 0.74 | 0.84 | 0.89 | 0.31 | 0.87 | 0.52 | 0.83 |
F1 | 0.42 | 0.78 | 0.84 | 0.89 | 0.31 | 0.87 | 0.52 | 0.85 | |
Total Accuracy | 0.27 | 0.52 | 0.58 | 0.63 | 0.40 | 0.81 | 0.42 | 0.65 | |
Total F1 | 0.28 | 0.59 | 0.64 | 0.71 | 0.43 | 0.83 | 0.45 | 0.71 |
Model | Mixtral | GPT-3.5-Turbo | GPT-4o | Total | |
Level | Metric | strict | flexible | strict | flexible | strict | flexible | strict | flexible |
1 | Accuracy | 0.25 | 0.41 | 0.34 | 0.42 | 0.53 | 0.54 | 0.37 | 0.46 |
F1 | 0.28 | 0.48 | 0.41 | 0.53 | 0.62 | 0.64 | 0.44 | 0.55 | |
2 | Accuracy | 0.08 | 0.30 | 0.43 | 0.50 | 0.54 | 0.72 | 0.35 | 0.51 |
F1 | 0.08 | 0.35 | 0.49 | 0.59 | 0.55 | 0.75 | 0.37 | 0.56 | |
3 | Accuracy | 0.09 | 0.35 | 0.28 | 0.40 | 0.59 | 0.74 | 0.32 | 0.50 |
F1 | 0.10 | 0.46 | 0.35 | 0.52 | 0.60 | 0.75 | 0.35 | 0.58 | |
4.1 | Accuracy | 0.15 | 0.47 | 0.77 | 0.77 | 0.71 | 0.76 | 0.54 | 0.67 |
F1 | 0.15 | 0.49 | 0.77 | 0.78 | 0.71 | 0.75 | 0.54 | 0.69 | |
4.2 | Accuracy | 0.00 | 0.47 | 0.75 | 0.76 | 0.73 | 0.74 | 0.49 | 0.66 |
F1 | 0.00 | 0.56 | 0.75 | 0.76 | 0.73 | 0.75 | 0.49 | 0.69 | |
Total Accuracy | 0.11 | 0.39 | 0.51| 0.57 | 0.62 | 0.70 | 0.42 | 0.56 | |
Total F1 | 0.12 | 0.47 | 0.56 | 0.65 | 0.64 | 0.73 | 0.44 | 0.61 |
Model | Mixtral | GPT-3.5-Turbo | GPT-4o | Total | |
Level | Metric | strict | flexible | strict | flexible | strict | flexible | strict | flexible |
1 | Accuracy | 0.33 | 0.43 | 0.32 | 0.41 | 0.51 | 0.54 | 0.39 | 0.46 |
F1 | 0.37 | 0.50 | 0.38 | 0.52 | 0.60 | 0.64 | 0.45 | 0.55 | |
2 | Accuracy | 0.14 | 0.35 | 0.39 | 0.46 | 0.28 | 0.72 | 0.27 | 0.51 |
F1 | 0.15 | 0.45 | 0.45 | 0.56 | 0.28 | 0.77 | 0.29 | 0.59 | |
3 | Accuracy | 0.12 | 0.39 | 0.38 | 0.45 | 0.09 | 0.66 | 0.20 | 0.50 |
F1 | 0.13 | 0.46 | 0.47 | 0.55 | 0.09 | 0.65 | 0.23 | 0.55 | |
4.1 | Accuracy | 0.07 | 0.47 | 0.71 | 0.75 | 0.55 | 0.76 | 0.44 | 0.66 |
F1 | 0.07 | 0.48 | 0.71 | 0.75 | 0.55 | 0.75 | 0.44 | 0.66 | |
4.2 | Accuracy | 0.31 | 0.56 | 0.70 | 0.74 | 0.18 | 0.73 | 0.40 | 0.68 |
F1 | 0.31 | 0.61 | 0.70 | 0.74 | 0.18 | 0.74 | 0.40 | 0.70 | |
Total Accuracy | 0.19 | 0.44 | 0.50 | 0.56 | 0.32 | 0.68 | 0.34 | 0.56 | |
Total F1 | 0.20 | 0.50 | 0.55 | 0.63 | 0.34 | 0.71 | 0.36 | 0.61 |
@@
Elaborate instructions without examples do not necessarily yield better results. Upon analyzing both types of prompts across all levels, it appears that augmenting the prompt with more information without examples does not consistently enhance performance. Level 3 prompts, when evaluated rigorously, exhibit an average decline of around in accuracy and F1 score compared to levels 1 and 2. When evaluated using more flexible metrics, the discrepancy diminishes to almost zero. GPT-4o tends to increase its performance with each level when using hand-written prompts, while it dramatically decreases it for the model-rephrased ones. The other two models consistently lower their metrics at the third level, especially Mixtral-8x7b, which can be attributed to the inclusion of explanatory text, as it strives to replicate the input text.
@@
ICL and COT prompting techniques lead to best results. All models scored their best results when prompted at levels 4.1 and 4.2, no matter the dataset or prompting template. Only GPT-4o had its best results for strict metrics at the first level when prompts where model-rephrased, which could be attributed to poor paraphrasing for the latter levels. It is no surprise that such models work best when an adequate output example is given, as literature [11] suggests. However, as Mixtral-8x7b sometimes provides explanations for its output, erroneous reasoning steps are noticeable, especially in cases where the input text contains a class type that is not present in the ontology. Thus, despite the GPT models exhibiting this behavior less frequently, LLMs still have significant room for improvement in terms of reasoning capabilities.
@@
Mixtral-8x7b rarely follows the required output format. The two metric measurement paradigms offer valuable insights into the models capacity to follow the given prompts. While GPT 3.5-turbo exhibits minimal disparity between the two perspectives, Mixtral 8x7b rarely produces texts that align with the specified template. Common errors include the addition of explanatory text (as evidenced by the 0 scores at the 4.2 level in table 5) or the full IRI of an entity. When strictly evaluated, the open-source model only tops accuracy on the Template Easy (TE), while on flexible paradigm it reaches accuracy on the same dataset. GPT-3.5 outputs accuracy under both metrics measurements, while the GPT-4o variant yields .
@@
Asking models to rephrase the system prompt might generally be a good idea for Mixtral-8x7b. Some experiments in the literature [10] ask the LLMs to formulate prompts for a given task. Inspired by it, we ask the LLMs to rephrase our manually written prompts to better align with their capabilities. As a comparison, Mixtral-8x7b benefits the most under rigorous evaluation, with an average increase of for each accuracy and F1 score. GPT-3.5-Turbo seems to conserve its behavior, signaling an increase of only . Surprisingly, GPT-4o exhibit a significant decrease in performance when it paraphrased the input prompts. On average, it lowered its performance by for both metrics, with third level prompts being the worst affected. Nonetheless, it’s promising to see the open-source model enhancing its output by closely adhering to the provided system prompt.